# Believing is Seeing: Unobserved Object Detection using Generative Models

### v2

@@ Appendix D Detailed methodology for VLM sampling
Visual-language models (VLMs) were queried using an automated pipeline between November 1st and 2nd, 2024. The latest official APIs were used for Claude 3.5 Sonnet [3] (claude-3-5-sonnet-20241022), ChatGPT-4o [47] (chatgpt-4o-latest), and Gemini 1.5 Pro [20] (gemini-1.5-pro-latest). LLaVa-34B-v1.6 (Nous-Hermes-2-34B) [37, 38], accessed via the Replicate API [^3] https://replicate.com/yorickvp/llava-v1.6-34b, is open source (unlike the other proprietary models) but was evaluated through APIs, as we could not perform local inference owing to model size.
@@ Appendix E Additional Results
In this section we provide all the complete ablation tables for individual datasets—NYU Depth V2 [46] and RealEstate10k [87]. We tabulate the results of the ablation of Diffusion with Forward Model samples [73] in 2D Tabs. 7 and 8, in 2.5D Tabs. 9 and 10 and in 3D Tabs. 11 and 12. We apply the same random seed to select 10 or 15 samples for random drop-outs, ensuring consistency in the sample selection process, retaining the points and the confidences from the input poses in all cases.
@@
We also provide the breakdown for the analysis of the SDXL-based analysis in 2D Tabs. 13 and 14 and the analysis in 2.5D Tabs. 16 and 15. In Tab. 3, we tabulated the combined ablation study using the total number of valid samples in both datasets. We present the following metrics along with their means and standard deviations: normalized entropy ($\mathcal{H}$), normalized cross-entropy ($\mathcal{H}^{\times}$), normalized nearest neighbor distance as a percentage ($\Delta$), 2D region-wise accuracy ($\mathcal{A}$), and false negative rates (FNR).
@@
SDXL w/ object cues $1.257\pm 2.033$ $0.848\pm 0.150$ $0.446\pm 25.490$ $0.0385$ $0.918\pm 0.147$
@@
SDXL w/ object cues $1.223\pm 2.107$ $0.778\pm 0.212$ $0.510\pm 26.929$ $0.0544$ $0.944\pm 0.142$
@@
Additional qualitative results.
In this section, we present additional qualitative results to supplement our analysis. Samples generated using the DFM are illustrated in Fig. 10, while those from SDXL are shown in Fig. 11. As expected, the DFM-based model exhibits poor qualitative performance on the NYU dataset due to its limited ability to generalize to out-of-distribution data [73]. The DFM-based model, as depicted in rows 6 and 7 of Fig. 12, frequently predicts flat distributions. While these distributions capture uncertainty and can localize the ground truth position, their higher uncertainty reflects in the normalization process followed by log-scaling, resulting in a wide probability mass spread across the distribution ($\mathcal{D}_{\mathcal{I}o}^{3D}$). For top-down heatmaps, we visualize the maximum normalized confidence value along the $+y$-axis for each column of the voxel grid. These visualizations may show spillover into regions without detected objects, which is an artifact of the softmax normalization and voxelization processes. In some 2D heatmaps, ground truth positions are omitted when the object is not visible in the ground truth image but is present in the scene. For SDXL-generated samples, we observe that semantic quality is significantly influenced by the guidance scale and the text prompt. When object cues are provided, both qualitative and quantitative performance improve noticeably. Conversely, generation without object cues or prompts yields minimal differences in quantitative results, as both scenarios result in limited detections. The qualitative results for SDXL analysis (column 3 in Figs. 5(d) and 12) are based on samples generated using prompts that include object cues. In certain SDXL-based heatmaps, spillovers into the central region may occur because the detector perceives the entire object, shared between the input image and the outpainted region, as part of the object.
(a) With object prompt: bed
(b) No object prompt
@@
(d) With object prompt: TV
(e) No object prompt
@@
(g) With object prompt: oven
(h) No object prompt
@@
(j) With object prompt: book
(k) No object prompt
@@
(m) With object prompt: oven
(n) No object prompt

### v3

@@ Appendix D Detailed methodology for VLM sampling
Visual-language models (VLMs) were queried using an automated pipeline between November 1st and 2nd, 2024. The latest official APIs were used for Claude 3.5 Sonnet [3] (claude-3-5-sonnet-20241022), ChatGPT-4o [46] (chatgpt-4o-latest), and Gemini 1.5 Pro [19] (gemini-1.5-pro-latest). LLaVa-34B-v1.6 (Nous-Hermes-2-34B) [37, 38], accessed via the Replicate API [^4] Replicate LLaVa-v1.6-34B model API documentation, is open source (unlike the other proprietary models) but was evaluated through APIs, as we could not perform local inference owing to model size.
@@ Appendix E Additional Results
In this section we provide all the complete ablation tables for individual datasets—NYU Depth V2 [65] and RealEstate10k [87]. We tabulate the results of the ablation of Diffusion with Forward Model samples [73] in 2D Tabs. 8 and 9, in 2.5D Tabs. 10 and 11 and in 3D Tabs. 12 and 13. We apply the same random seed to select 10 or 15 samples for random drop-outs, ensuring consistency in the sample selection process, retaining the points and the confidences from the input poses in all cases.
@@
We also provide the breakdown for the analysis of the SDXL-based analysis in 2D Tabs. 14 and 15 and the analysis in 2.5D Tabs. 17 and 16. In Tab. 3, we tabulated the combined ablation study using the total number of valid samples in both datasets. We present the following metrics along with their means and standard deviations: normalized entropy ($\mathcal{H}$), normalized cross-entropy ($\mathcal{H}^{\times}$), normalized nearest neighbor distance as a percentage ($\Delta$), 2D region-wise accuracy ($\mathcal{A}$), and false negative rates (FNR).
Additional qualitative results.
In this section, we present additional qualitative results to supplement our analysis. Samples generated using the DFM are illustrated in Fig. 10, while those from SDXL are shown in Fig. 12. As expected, the DFM-based model exhibits poor qualitative performance on the NYU dataset due to its limited ability to generalize to out-of-distribution data [73]. The DFM-based model, as depicted in rows 6 and 7 of Fig. 13, frequently predicts flat distributions. While these distributions capture uncertainty and can localize the ground truth position, their higher uncertainty reflects in the normalization process followed by log-scaling, resulting in a wide probability mass spread across the distribution ($\mathcal{D}_{\mathcal{I}o}^{3D}$). For top-down heatmaps, we visualize the maximum normalized confidence value along the $+y$-axis for each column of the voxel grid. These visualizations may show spillover into regions without detected objects, which is an artifact of the softmax normalization and voxelization processes. In some 2D heatmaps, ground truth positions are omitted when the object is not visible in the ground truth image but is present in the scene. For SDXL-generated samples, we observe that semantic quality is significantly influenced by the guidance scale and the text prompt. When object cues are provided, both qualitative and quantitative performance improve noticeably. Conversely, generation without object cues or prompts yields minimal differences in quantitative results, as both scenarios result in limited detections. The qualitative results for SDXL analysis (column 3 in Figs. 5(d) and 13) are based on samples generated using prompts that include object cues. In certain SDXL-based heatmaps, spillovers into the central region may occur because the detector perceives the entire object, shared between the input image and the outpainted region, as part of the object.
Failure case analysis.
In our current DFM experiments, 57% of failures arise from object prediction errors (non-occluded), while 43% result from detector false negatives in ambiguous contexts (e.g., multiple doors in a corridor). No failures stem from the inability to predict occluded objects. SDXL-based methods with object prompts show no failures. In VLM experiments, all failures are direct prediction errors of the model.
Outdoor Scenes.
Our framework is also applicable to outdoor scenes. We demonstrate this with an outdoor scene from the RealEstate10k dataset featuring the COCO object ‘car’ (Fig. 11(d), Tab. 7). The SDXL and VLM-based models perform well as expected while the DFM struggles in outdoor settings due to limited training on large-scale outdoor data, retraining which exceeds our computational resources.
Table 7: Results for the outdoor scene–object pair in Fig. 11(d).
Methods 2D Experiment 2.5D Experiment 3D Experiment
$\mathcal{H}^{\times}\downarrow$ $\mathcal{H}\downarrow$ $\Delta(\%)\downarrow$ $\mathcal{A}\uparrow$$\mathcal{H}^{\times}\downarrow$ $\mathcal{H}\downarrow$ $\Delta(\%)\downarrow$$\mathcal{H}^{\times}\downarrow$ $\mathcal{H}\downarrow$ $\Delta(\%)\downarrow$
DFM (25 samples, 4 poses) 1.000 1.000 $\infty$ 0.670 0.998 1.000 $\infty$ 0.998 1.000 $\infty$
SDXL (w. obj prompt) 0.891 0.771 0.000 0.670 0.788 0.962 0.000
ChatGPT-4o 0.880 0.879 0.000 0.670
Claude 3.5 Sonnet 0.880 0.879 0.000 0.670
Gemini 1.5 Pro 0.994 0.993 0.000 0.670
LLaVa-v1.6-34b 0.999 0.998 0.000 0.670
@@
SDXL w/ object cues $1.257\pm 2.033$ $0.848\pm 0.150$ $0.446\pm 25.490$ $0.039$ $0.918\pm 0.147$
@@
SDXL w/ object cues $1.223\pm 2.107$ $0.778\pm 0.212$ $0.510\pm 26.929$ $0.054$ $0.944\pm 0.142$
@@
(a) No object prompt
(b) With object prompt: bed
@@
(d) No object prompt
(e) With object prompt: TV
@@
(g) No object prompt
(h) With object prompt: oven
@@
(j) No object prompt
(k) With object prompt: book
@@
(m) No object prompt
(n) With object prompt: oven